### v3

@@ 4.2 EHR Text Serialization
To use LLM embedding models for representing EHR records, we serialized the records into textual formats. We limited the serialization length to 4,096 tokens (approximately 16,000 characters) due to computational constraints (see Section 4.6). Two experimental runs were conducted with serializations extending up to 8,192 tokens. The primary goal was to create a detailed and informative serialization requiring minimal preprocessing while ensuring that medically relevant information appeared early in the text. This approach mitigated truncation risks in lengthy records, preserving critical details even when older entries were omitted. To convert the visits and clinical events in the EHRSHOT dataset into text, we used the semantic information embedded in the dataset. Each clinical event was labeled using the format “ontology/code”. EHRSHOT provided a set of prepared ontologies for resolving concept codes into their descriptions, which we incorporated. An analysis was performed to evaluate the use of these ontologies for all events of a subset of 200 patients across the task groups for operational outcomes and new diagnoses, covering 2,968 labels. We identified the following ontologies: Logical Observation Identifiers Names and Codes, SNOMED, RxNorm, CPT4, Domain, CARE_SITE, RxNorm Extension, Medicare Specialty, ICD10PCS, CMS Place of Service, Cancer Modifier, ICD9Proc, CVX, ICDO3, HCPCS, OMOP Extension, Condition Type. We excluded ontologies containing only a single value (Domain, Medicare Specialty, CMS Place of Service, OMOP Extension, and Condition Type). Codes of the ontologies CPT4, CARE_SITE, ICD10PCS, Cancer Modifier, CVX, and ICDO3 were not resolved with the provided ontologies. Cancer Modifier codes contained UICC cancer stages that we parsed manually. For CPT4, ICD10PCS, and CVX we used custom mapping files that we manually added.333We downloaded CPT4 from https://gist.github.com/lieldulev/439793dc3c5a6613b661c33d71fdd185, ICD10PCS from https://hcup-us.ahrq.gov/toolssoftware/procedureicd10/procedure_icd10_archive.jsp, and CVX from https://www2a.cdc.gov/vaccines/iis/iisstandards/vaccines.asp?rpt=cvx. We excluded CARE_SITE and ICDO3 since we found no way to resolve these to useful descriptions.
@@
Various approaches exist for serializing structured data, including row-by-row serialization [42], template-based methods [43], or structured data formats like JSON, HTML, and Markdown [44]. We used Markdown due to its minimal overhead and overall benefits of using a structured data input format for LLMs [45]. To harmonize dates, all timestamps were normalized relative to January 1, 2024, designated as the prediction reference time. Serialization began with patient demographics, typically the first event for each patient, where birthdates were converted into ages (in years) for simplicity. Since 65% of the dataset comprised time-series data encoded via LOINC, which we found imbalanced, we aggregated LOINC-coded events. Using the same patient subset as the ontology analysis, we identified the most frequent codes and categorized them into vital signs, body metrics, and lab values, selecting 24 key medical concepts. To avoid duplicates, we merged synonymous codes (see Table 7). The last three values of each concept are presented, and we filtered implausible values. To further enrich the text representation, we manually added default units and assessments (low, normal, high) based on standard ranges (see Table 7). Following the aggregated data, a summary of all visits was included to address the potential truncation of older visits. Events not associated with visits were then presented, using the same aggregation logic to display the last three values where applicable. Finally, a detailed chronological presentation of all visits was included, with events categorized into conditions (SNOMED, Visit, Cancer Modifier, CVX, HCPCS), medications (RxNorm, RxNorm Extension), and procedures (CPT4, ICD10PCS, ICD9Proc).
4.3 Potential Bias of Manually Defining an EHR Text Serialization
Defining an EHR serialization involved subjective decisions, which may have introduced bias. For instance, awareness of the prediction tasks could influence the prioritization of certain data elements, potentially favoring task-relevant information. To minimize this risk, we aimed to create an objectively defined serialization that encapsulates key aspects of the EHR records. Also, due to computational constraints, we evaluated only three alternatives to our final serialization for 4,096 tokens: (1) appending a list of all unique conditions at the beginning, (2) omitting the three values for the listed comments (e.g., for “Cigarette consumption” and “pH measurement, venous” in Fig. 2), and (3) combining both approaches. For GTE-Qwen2-7B and LLM2Vec-Llama-3.1-8B, the EHR serialization chosen for our experiments performed on par or better than the three alternatives. We also selected the current serialization for its simplicity and completeness of medical information. This approach avoided introducing additional entries of all unique conditions and preserved the most recent values for visit-level concepts.
@@ 4.4 LLM Embedding Models and Baselines
In this study, we evaluated two LLM embedding models, GTE-Qwen2-7B and LLM2Vec-Llama-3.1-8B, based on state-of-the-art decoder-only LLMs. We selected these models for their ability to handle the 4,096-token EHR serializations used in our experiments. For comparison, we also tested a smaller variant of both models. As additional baselines, we included commonly used encoder-only embedding models with smaller input sizes (512 tokens). To use them with 4,096 token inputs, the EHR serializations were split into up to eight 512-token chunks, and the resulting embeddings were averaged to generate a single representation. For all LLM embedding and smaller language models, we used mean pooling of the last layer for the final embedding [27, 46]. The LLM2Vec models used a slight variation that only incorporates the tokens that do not belong to the instruction. Below is an overview of all models:
GTE-Qwen2-7B
@@
This LLM embedding model is based on the Qwen2-7B-Instruct LLM [32], which uses a decoder-only Transformer architecture with 28 layers, 28 attention heads, and a hidden size of 3,584444Hugging Face identifier: Alibaba-NLP/gte-Qwen2-7B-instruct. It was trained with autoregressive next token prediction and converted into an embedding model using the General Text Embedding (GTE) method [30]. This conversion replaces causal attention with bidirectional attention, enabling the model to attend to both left and right contexts for token embedding, akin to BERT. Contrastive learning was applied using a mixture of private datasets to enhance embedding performance. The model also incorporates instructions tailored for embedding tasks and supports a context size of up to 32,000 tokens.
GTE-Qwen2-1.5B
A smaller variant of GTE-Qwen2-7B, this model is based on Qwen2-1.5B-Instruct, with 28 layers, 12 attention heads, and a hidden size of 1,536. It was also trained using the GTE method [30] and supports a context size of up to 32,000 tokens.
LLM2Vec-Llama-3.1-8B
@@
This model is built upon the Llama-3.1-8B-Instruct LLM [33], which has a decoder-only Transformer architecture with 32 layers, 32 attention heads, and a hidden size of 4,096555Hugging Face identifier: McGill-NLP/LLM2Vec-Meta-Llama-31-8B-Instruct-mntp-supervised. Initially trained for next-token prediction, it was converted to an embedding model using the LLM2Vec method [27]. This method adds bidirectional attention and fine-tunes the model with supervised contrastive learning on embedding tasks. The fine-tuning used curated data from the public E5 dataset [47, 48], containing approximately 1.5 million entries. The model supports task-specific instructions and supports a context size of up to 128,000 tokens.
LLM2Vec-Llama-2-1.3B
@@
DeBERTa v3 is an encoder-only Transformer model designed for token embeddings [50]. It improves upon its predecessor by replacing the masked language modeling objective with replaced token detection and using Gradient-Disentangled Embedding Sharing. We evaluated the base variant (12 layers, 12 attention heads, 768 hidden size) and the large variant (24 layers, 12 attention heads, 1,024 hidden size), with parameter counts of 183M and 434M, respectively666Hugging Face identifiers: microsoft/deberta-v3-base,large.
@@
BERT is a well-established text embedding model using an encoder-only Transformer trained with the masked language modeling objective [15]. We included both the base (12 layers, 12 attention heads, 768 hidden size, 110M parameters) and large (24 layers, 16 attention heads, 1,024 hidden size, 340M parameters) variants as benchmarks777Hugging Face identifiers: google-bert/bert-base,large-uncased. While not state-of-the-art, BERT models remain widely used in embedding tasks.
@@
CLMBR-T-Base is a specialized EHR foundation model trained on 2.57 million de-identified EHRs from Stanford Medicine with autoregressive next code prediction [13, 31]. It uses gated recurrent units and has 12 layers and a hidden dimension of 768999Hugging Face identifier: StanfordShahLab/clmbr-t-base. The model has 141M parameters and allows for a context window of 496 codes. CLMBR-T-Base has demonstrated consistent improvements over count-based baselines for a variety of clinical prediction tasks [31]. It serves as a main baseline for our experiments to test specialized EHR models against general-purpose text embedding models for representing EHR records.
@@
To test whether the LLM embedding models and the EHR foundation model learn orthogonal information, we combined both models for the prediction. To this end, we simply appended both embeddings. The resulting embeddings have dimensions 4,352 (GTE-Qwen2-7B) and 4,864 (LLM2Vec-Llama-3.1-8B).
@@
Count models have proven to be strong baselines for EHR prediction tasks [6, 12, 13]. The basic idea is to encode all EHR events of a patient in a single vector where each entry represents the number of occurrences of a medical concept. We used the counts baseline introduced in [31] that further extends this approach with ontology expansion, enriching the vectors with parent and child concepts.
Based on the embeddings or the counts vectors generated by the methods described above, a classification head was trained and validated for each prediction task. For the embedding models we used a logistic regression head. For the count-based model, we used a GBM [41], which proved superior [31]. We adopted the parameter tuning of the classification heads from the EHRSHOT benchmark to ensure comparability of results.
@@ 4.5 Instructions for LLM Embedding Models
The GTE and LLM2Vec models used instruction-tuned embeddings, requiring task-specific prompts. Hence, we added simple instructions for each prediction task based on their respective instruction templates. For instance, for prediction of anemia, we added “Given a patient’s electronic healthcare record (EHR) in Markdown format, retrieve relevant passages that answer the query: has the patient anemia”. The existing EHRSHOT benchmark encoded the EHRs of the same patient and the identical prediction times only once for efficiency reasons. However, to support task-specific instructions, we had to change this default behavior, leading to 1,161,412 instead of 406,379 EHRs to encode, resulting in longer processing times. The difference between 1,161,412 labels used in our experiments and the total number of labels of 1,178,665 (Table 2) is because some labels share the same task and prediction time and are therefore merged. For assignment of new diagnoses tasks in UKB, the same prompt was used for all diseases to reduce computational resources. We list all instructions in Table 8 and perform ablations to test the effect of the instructions.

### v4

@@ 4.2 EHR Text Serialization
To use LLM embedding models for representing EHR records, we serialized the records into a textual format agnostic to the prediction task. Following the EHRSHOT benchmark, we added all events before the label time. We limited the serialization length to 8,192 tokens (approximately 32,000 characters) due to computational constraints (see Section 4.8). Some additional experiments were conducted with a 4,096-token limit due to computational constraints. The primary goal was to create a detailed and informative serialization requiring minimal preprocessing while ensuring that medically relevant information appeared early in the text. This approach mitigated truncation risks in lengthy records, preserving critical details even when older entries were omitted. To convert the visits and clinical events in the EHRSHOT dataset into text, we used the semantic information embedded in the dataset. Each clinical event was labeled using the format “ontology/code”. EHRSHOT provided a set of prepared ontologies for resolving concept codes into their descriptions, which we incorporated. An analysis was performed to evaluate the use of these ontologies for all events of a subset of 200 patients across the task groups for operational outcomes and new diagnoses, covering 2,968 labels. We identified the following ontologies: Logical Observation Identifiers Names and Codes, SNOMED, RxNorm, CPT4, Domain, CARE_SITE, RxNorm Extension, Medicare Specialty, ICD10PCS, CMS Place of Service, Cancer Modifier, ICD9Proc, CVX, ICDO3, HCPCS, OMOP Extension, Condition Type. We excluded ontologies containing only a single value (Domain, Medicare Specialty, CMS Place of Service, OMOP Extension, and Condition Type). Codes of the ontologies CPT4, CARE_SITE, ICD10PCS, Cancer Modifier, CVX, and ICDO3 were not resolved with the provided ontologies. Cancer Modifier codes contained UICC cancer stages that we parsed manually. For CPT4, ICD10PCS, and CVX we used custom mapping files that we manually added.333We downloaded CPT4 from https://gist.github.com/lieldulev/439793dc3c5a6613b661c33d71fdd185, ICD10PCS from https://hcup-us.ahrq.gov/toolssoftware/procedureicd10/procedure_icd10_archive.jsp, and CVX from https://www2a.cdc.gov/vaccines/iis/iisstandards/vaccines.asp?rpt=cvx. We excluded CARE_SITE and ICDO3 since we found no way to resolve these to useful descriptions.
@@
Various approaches exist for serializing structured data, including row-by-row serialization [54], template-based methods [55], or structured data formats like JSON, HTML, and Markdown [56]. We used Markdown due to its minimal overhead and overall benefits of using a structured data input format for LLMs [57]. To standardize temporal information, all dates were normalized relative to January 1, 2024, the prediction reference time. Only past visit dates were included, as they were considered most relevant and helped keep the serialization concise. To highlight temporal proximity, each date was annotated with the number of days before the prediction time. Serialization began with patient demographics, typically the first event for each patient, where birthdates were converted into ages (in years) for simplicity. Since 65% of the dataset comprised time-series data encoded via LOINC, which we found to be imbalanced, we aggregated LOINC-coded events. Using the same patient subset as the ontology analysis, we identified the most frequent codes and categorized them into vital signs, body metrics, and lab values, selecting 24 key medical concepts. To avoid duplicates, we merged synonymous codes (see Table 8). The last three values of each concept are presented, and we filtered implausible values. To further enrich the text representation, we manually added default units and assessments (low, normal, high) based on standard ranges (see Table 8). Following the aggregated data, a summary of all visits was included to address the potential truncation of older visits. Events not associated with visits were then presented, using the same aggregation logic to display the last three values where applicable. Finally, a detailed reverse-chronological presentation of all visits was included, with events categorized into conditions (SNOMED, Visit, Cancer Modifier, CVX, HCPCS), medications (RxNorm, RxNorm Extension), and procedures (CPT4, ICD10PCS, ICD9Proc).
4.3 Testing Alternative EHR Text Serializations
Our EHR text serialization contained several design choices and may lack generalizability. Specifically, we knew the EHRSHOT prediction tasks before designing the serialization, which could have introduced bias in the design of the EHR serialization. In total, we initially evaluated three targeted alternatives with a 4,096-token limit due to computational constraints: (1) appending a list of all unique conditions at the beginning, (2) omitting the three values for the listed concepts (e.g., for “Cigarette consumption” and “pH measurement, venous” in Fig. 2), and (3) combining both approaches. We chose the current serialization due to initial experiments with Qwen2-Emb-7B and Llama3.1-LLM2Vec-8B, where it performed on par or better than the three alternatives. To test the effect of our serialization format more broadly, we also evaluated a very simple serialization by creating a newline-separated list of all medical events of a patient ordered by time. We tested four variants with different sorting directions and with and without timestamps.
Additionally, we tested three data formats by converting the structured Markdown serialization into JSON, XML, and YAML. We used standard Python libraries for the conversion process. Due to different formatting overheads of the formats, the amount of clinical content included under a fixed token budget may differ.
@@ 4.4 LLM Embedding Models and Baselines
We evaluated three LLM embedding models: Qwen3-Embedding-8B (Qwen3-Emb-8B) [38, 43], GTE-Qwen2-7B-Instruct (Qwen2-Emb-7B) [37, 42], and LLM2Vec-Llama-3.1-8B-Instruct (Llama3.1-LLM2Vec-8B) [48, 39], based on state-of-the-art decoder-only LLMs. We selected these models for their ability to handle the 8,192-token EHR serializations used in our experiments. For comparison, we also tested smaller variants of Qwen3-Emb-8B and Qwen2-Emb-7B. All LLM-based models received task-specific instructions prepended to the EHR serialization (see Section 4.5), and we used each model’s default embedding configuration. For Qwen-based embedding models, we used the last-layer hidden state of the final ([EOS]) token as the patient embedding [37, 43]. For LLM2Vec models, we applied mean pooling over the last-layer hidden states corresponding to the EHR serialization, excluding the instruction tokens from pooling [39]. Importantly, the instruction still conditions the internal representation and thus influences the resulting embedding.
@@
As additional baselines, we included commonly used encoder-only embedding models with shorter input limits (512 tokens). To apply them to 8,192-token inputs, we split each input into up to 16 chunks of 512 tokens and averaged the chunk embeddings to obtain a single representation. These encoder-only models did not use task instructions, and embeddings were computed via mean pooling over the last hidden layer. Below is an overview of all models.
Qwen3-Embedding-8B/4B/0.6B
This model family is built on the Qwen3 foundation LLMs [38] and trained as instruction-aware embedding models via a multi-stage recipe [43]. All variants use decoder-only Transformers with causal attention and produce the embedding from the last-layer hidden state of the final [EOS] token (see Section 4.5). The 8B and 4B models have 36 layers, while the 0.6B model has 28 layers; all support a context window of up to 32,000 tokens. By default, the embedding dimensionality is 4,096 (8B), 2,560 (4B), and 1,024 (0.6B), with support for flexible output dimensions. Training comprises large-scale weakly supervised pretraining on synthetic query–document pairs followed by supervised fine-tuning on labeled and filtered synthetic data optimizing a contrastive learning objective. Finally, different checkpoints are merged to improve robustness [43].
Qwen2-Embedding-7B/1.5B
@@
Built on GTE-Qwen2-7B-Instruct and GTE-Qwen2-1.5B-Instruct [37], these models use decoder-only Transformers (7B: 28 layers, 28 heads, hidden size 3,584; 1.5B: 28 layers, 12 heads, hidden size 1,536). They are first trained with autoregressive next-token prediction and then converted to embedding models via the General Text Embedding (GTE) procedure [42], which replaces the causal mask with a BERT-like bidirectional attention during embedding extraction. Contrastive learning was applied using a mixture of private datasets to enhance embedding performance. The models also incorporate instructions tailored for embedding tasks and support a context size of up to 32,000 tokens.
Llama3.1-LLM2Vec-8B
@@
This model is built upon the Llama-3.1-8B-Instruct LLM [48], which has a decoder-only Transformer architecture with 32 layers, 32 attention heads, and a hidden size of 4,096444Hugging Face identifier: McGill-NLP/LLM2Vec-Meta-Llama-31-8B-Instruct-mntp-supervised. Initially trained for next-token prediction, it was converted to an embedding model using the LLM2Vec method [39]. This method adds bidirectional attention and fine-tunes the model with supervised contrastive learning on embedding tasks. The fine-tuning used curated data from the public E5 dataset [58, 59], containing approximately 1.5 million entries. The model supports task-specific instructions and supports a context size of up to 128,000 tokens.
@@
DeBERTa v3 is an encoder-only Transformer model designed for token embeddings [60]. It improves upon its predecessor by replacing the masked language modeling objective with replaced token detection and using Gradient-Disentangled Embedding Sharing. We evaluated the base variant (12 layers, 12 attention heads, 768 hidden size) and the large variant (24 layers, 12 attention heads, 1,024 hidden size), with parameter counts of 183M and 434M, respectively555Hugging Face identifiers: microsoft/deberta-v3-base,large.
@@
BERT is a well-established text embedding model using an encoder-only Transformer trained with the masked language modeling objective [15]. We included both the base (12 layers, 12 attention heads, 768 hidden size, 110M parameters) and large (24 layers, 16 attention heads, 1,024 hidden size, 340M parameters) variants as benchmarks666Hugging Face identifiers: google-bert/bert-base,large-uncased. While not state-of-the-art, BERT models remain widely used in embedding tasks.
@@
MedBERT
This model builds on BioClinicalBERT via continued domain pretraining on heterogeneous biomedical corpora (N2C2, BioNLP, CRAFT, and biomedical Wikipedia) and was originally introduced for biomedical NER [63]. We include it as a domain-adapted baseline for comparison888Hugging Face identifier: Charangan/MedBERT.
@@
CLMBR-T-Base is a specialized EHR foundation model trained on 2.57 million de-identified EHRs from Stanford Medicine with autoregressive next code prediction [13, 44]. It uses transformers and has 12 layers and a hidden dimension of 768999Hugging Face identifier: StanfordShahLab/clmbr-t-base. The model has 141M parameters and allows for a context window of 496 codes. CLMBR-T-Base has demonstrated consistent improvements over count-based baselines for a variety of clinical prediction tasks [44]. It serves as a main baseline for our experiments to test specialized EHR models against general-purpose text embedding models for representing EHR records.
@@
To test whether the LLM embedding models and the EHR foundation model learn orthogonal information, we combined both models for the prediction. To this end, we simply appended both embeddings. The resulting embeddings have dimensions 4,864 for Qwen3-Emb-8B.
@@
Count models have proven to be strong baselines for EHR prediction tasks [6, 12, 13].
The basic idea is to encode all EHR events of a patient in a single vector where each entry represents the number of occurrences of a medical concept.
We used the count-based baseline introduced in [44], which uses counts from different time intervals and extends this approach with ontology expansion, enriching the vectors with parent and child concepts.
Notably, this baseline does not incorporate any values associated with medical concepts, such as laboratory measurements, vital signs, or other continuous fields, but only the occurrence counts of each concept code.
Based on the embeddings or the counts vectors generated by the methods described above, a classification head was trained and validated for each prediction task.
For the embedding models we used a logistic regression head. For the count-based model, we used a GBM [53], which proved superior [44].
We adopted the parameter tuning of the classification heads from the EHRSHOT benchmark to ensure comparability of results.
@@ 4.5 Instructions for LLM Embedding Models
The Qwen and LLM2Vec models used instruction-tuned embeddings, requiring task-specific prompts. Hence, we added simple instructions for each prediction task based on their respective instruction templates. For instance, for prediction of anemia, we added “Given a patient’s electronic healthcare record (EHR) in Markdown format, retrieve relevant passages that answer the query: has the patient anemia”. The existing EHRSHOT benchmark encoded the EHRs of the same patient and the identical prediction times only once for efficiency reasons. However, to support task-specific instructions, we changed this default behavior and encoded each (patient, task, prediction time) instance, leading to 1,161,412 instead of 406,379 EHRs to encode and resulting in longer processing times. The difference between 1,161,412 labels used in our experiments and the total number of labels of 1,178,665 (see Table 2) is because some labels share the same task and prediction time and are therefore merged. Encoder-only language models did not use instructions and thus required encoding only the 406,379 EHRs from the original implementation, yielding a substantial computational advantage. For assignment of new diagnoses tasks in UKB, the same prompt was used for all diseases to reduce computational resources. We list all instructions in Table 9 and perform ablations to test the effect of the instructions.
4.6 Fine-Tuned LLM Embedding and LLM Decoder Models
We evaluated two approaches for clinical prediction: LLM embedding models that encode the EHR serialization into a fixed-length vector prior to classification, and LLM decoder models that generate text whose token probabilities can be used for classification. We further assessed the effect of fine-tuning for both approaches.
To this end, we implemented a pipeline for fine-tuning Qwen3-Emb-8B [38, 43] with LoRA adapters [64] applied to the encoder attention projections and MLP blocks. The encoder is wrapped in a lightweight classification head consisting of a dropout layer (dropout rate ) and a linear projection. Inputs are constructed by concatenating the task instruction and the EHR serialization, tokenized with enforced left padding to preserve relative recency, and batched with a custom collator that truncates to 4,096 tokens. The entire model is optimized end-to-end with cross-entropy loss. For comparability, both the encoder and decoder variants use the same LoRA hyperparameters (, , dropout ). Training uses the Hugging Face Trainer with cosine learning-rate decay, early stopping (patience five epochs), gradient checkpointing, and gradient accumulation chosen so that all variants share an effective batch size of . The models are trained with mixed precision (BF16), and zero-shot runs () are evaluated without parameter updates to the backbone.
The decoder pipeline adapts the instruction-tuned causal Qwen3-8B101010Hugging Face identifier: Qwen/Qwen3-8B checkpoint [38] with LoRA adapters applied to all attention projections and MLP up/down projections (q_proj, k_proj, v_proj, o_proj, up_proj, down_proj, gate_proj). Serialized visits are converted into chat-formatted prompts that additional prepend task instructions enforcing the output of Yes or No tokens (see Table 9). During LoRA fine-tuning, the model is trained to predict the literal tokens Yes or No as target labels. During scoring, the model predicts token probabilities for Yes and No at the final decoding step We aggregate probabilities across all variants of these tokens (including different capitalizations and punctuation) to yield calibrated decision scores without additional heads.
For both model families, evaluation computes AUROC, AUPRC, and Brier score on the held-out test set, alongside 1,000 patient-level bootstrap replicates to estimate standard deviations and 95% confidence intervals. To keep computational cost manageable, we used a token limit of 4,096 and evaluated the Anticipating Lab Test Results group on a uniformly random subset of 4,096 test samples in each run.
4.7 Existing Methods Using Language Models for EHR Prediction
Two broad approaches have emerged for leveraging language models with structured EHR data: using model generations directly for prediction and using vector embeddings derived from the model’s hidden states. Early work on the generation-based approach converted structured insurance claims into short text snippets and queried an encoder–decoder model with a 512-token context limit, reporting improvements over logistic regression in few-shot regimes (16–64 examples) for end-of-life, surgery, and length-of-stay tasks [26]. Subsequent studies explored more capable LLMs, prompt design, and fine-tuning strategies on serialized EHR inputs, generally treating prediction as text generation or calibrated scoring from logits [27, 28, 29, 30, 31, 32]. One variant constrains the model’s output space to medical codes to better align free-text reasoning with code-based clinical labels, thereby bridging the gap between general-purpose language models and EHR-specific foundation models [65].
In contrast to generation, embedding-based methods treat an EHR-to-text serialization as input and extract a fixed-dimensional representation for a downstream classifier [33, 34, 35]. Gao et al. provide a systematic comparison across serialization formats, LLMs, and embedding strategies on MIMIC-III and a private 660-patient clinical-deterioration dataset, and report that LLM embeddings paired with a prediction head can be competitive in some settings. However, raw numerical features remain strong baselines [33]. They also evaluate an encoder based approach producing Yes and No tokens in the zero-shot setting but find no discriminatory abilities for Mistral-7B-Instruct and Llama3-8B-Instruct. The DeLLiriuM method introduced in [35] also evaluates embeddings derived from different LLMs for delirium prediction and performs detailed model introspection via a SHAP analysis. Most similar to our work is the Multiple Embedding Model for EHR (MEME) method, which encodes different EHR modalities such as vitals, medications, and diagnoses into separate embeddings and fuses them with a self-attention layer [34]. This framework uses the relatively small and efficient MedBERT language model for embedding creation and reports superior performance to GPT-4 [34].
Our study follows the embedding paradigm but differs from prior work in three ways. First, we use LLMs explicitly adapted for embedding via contrastive learning. Second, we operate with substantially longer inputs (up to 8,192 tokens), which is important for longitudinal EHR serialization where earlier studies often processed up to 3,076 tokens [33] or relied on chunking [34]. Third, we evaluate on large public longitudinal cohorts and standardized benchmarks from EHRSHOT and UKB. We also include a MEME-style baseline with a linear head in place of self-attention due to data sparsity [34]. Finally, consistent with [33], we explicitly compare embeddings to direct LLM outputs and examine parameter-efficient fine-tuning to assess robustness and practical trade-offs (see Section 2.6).